### v1

# Dual-VQE: A quantum algorithm to lower bound the ground-state energy

Abstract
The variational quantum eigensolver (VQE) is a hybrid quantum–classical variational algorithm that produces an upper-bound estimate of the ground-state energy of a Hamiltonian. As quantum computers become more powerful and go beyond the reach of classical brute-force simulation, it is important to assess the quality of solutions produced by them. Here we propose a dual variational quantum eigensolver (dual-VQE) that produces a lower-bound estimate of the ground-state energy. As such, VQE and dual-VQE can serve as quality checks on their solutions; in the ideal case, the VQE upper bound and the dual-VQE lower bound form an interval containing the true optimal value of the ground-state energy. The idea behind dual-VQE is to employ semi-definite programming duality to rewrite the ground-state optimization problem as a constrained maximization problem, which itself can be bounded from below by an unconstrained optimization problem to be solved by a variational quantum algorithm. When using a convex combination ansatz in conjunction with a classical generative model, the quantum computational resources needed to evaluate the objective function of dual-VQE are no greater than those needed for that of VQE. We simulated the performance of dual-VQE on the transverse-field Ising model, and found that, for the example considered, while dual-VQE training is slower and noisier than VQE, it approaches the true value with error of order $10^{-2}$.
variational quantum eigensolver; semi-definite programming; slack variables; ground-state energy
preprint:
Introduction—A key approach hoped to achieve quantum advantage on near-term quantum computers involves the use of variational quantum algorithms [5, 14]. The variational quantum eigensolver (VQE) was the first such algorithm proposed [36]: it estimates an upper bound on the ground-state energy of a Hamiltonian. The basic idea behind it is simple, based on the well known variational principle from physics [22]. For a Hamiltonian $H$, its ground-state energy is equal to its minimum eigenvalue $\lambda_{\min}(H)$, which is equal to the following minimization problem:
$\lambda_{\min}(H)=\inf_{\rho\in\mathcal{D}}\operatorname{Tr}[H\rho]$ (1)
where $\mathcal{D}$ is the set of density matrices (positive semi-definite matrices with trace equal to one). Since the objective function $\operatorname{Tr}[H\rho]$ is linear in $\rho$ and the set of density matrices is convex, the optimal value of (1) is obtained on the boundary of the set $\mathcal{D}$ (i.e., the set of pure states). It then follows that
$\lambda_{\min}(H)=\inf_{|\psi\rangle\in\mathcal{P}}\langle\psi|H|\psi\rangle,$ (2)
where $\mathcal{P}$ is the set of state vectors. In a variational approach, the set of pure states is approximated by a parameterized set $\{|\psi(\theta)\rangle\}_{\theta\in\Theta}$ of states, where $\Theta$ is a parameter set and $\theta$ is a parameter vector. The equality in (2) is preserved if the set of parameterized pure states contains an optimal pure state. In reality, guaranteeing this condition is not possible, leading to the following upper bound on the ground-state energy:
$\lambda_{\min}(H)\leq\inf_{\theta\in\Theta}\langle\psi(\theta)|H|\psi(\theta)\rangle.$ (3)
An example of wide interest is when $n\in\mathbb{N}$ and $H$ is a $2^{n}\times 2^{n}$ Hermitian matrix, describing the local interactions that take place between $n$ spin-1/2 particles. In this case, $H=\sum_{j=1}^{L}H_{j}$, where each $H_{j}$ is a Hermitian matrix of constant size that acts non-trivially on a constant number of particles. Observing that the optimization in (1) is a semi-definite program (SDP) [40], one could attempt to solve it by means of well known approaches used to solve SDPs; however, the computational complexity of such approaches is exponential in $n$ and thus infeasible. Standard classical approaches for solving (1) in the general case suffer from the same problem, or they avoid it by not allowing for the optimization to include highly entangled states (see [18] for a review of some of these methods).
VQE instead parameterizes the set of state vectors by a parameterized quantum circuit $U(\theta)$ acting on an easily preparable initial state $|\psi_{0}\rangle$, leading to $|\psi(\theta)\rangle\coloneqq U(\theta)|\psi_{0}\rangle$, and measures the expectation of each observable $H_{j}$ with respect to $|\psi(\theta)\rangle$ in order to obtain an estimate $\widehat{E_{j}}$ of $\langle\psi(\theta)|H_{j}|\psi(\theta)\rangle$. Then the sum $\sum_{j}\widehat{E_{j}}$ is an estimate of the desired expectation $\langle\psi(\theta)|H|\psi(\theta)\rangle$. The estimates of $\langle\psi(\theta)|H|\psi(\theta)\rangle$ and its gradient can be fed to a classical optimization procedure and this process iterated until convergence or a maximum number of iterations is reached. As a consequence of (3), this procedure in the ideal case produces an estimate of an upper bound for the ground-state energy $\lambda_{\min}(H)$.
In the short term, while classical computers are still competitive with small-scale quantum computers, it is possible to test how well VQE is performing by comparing its value with an extremely precise estimate of the true known value. As mentioned above, an SDP can output such a precise estimate for sufficiently small system sizes. However, this approach is not viable in the long term, which leads to the key question that our paper addresses:
How can we assess the quality of VQE’s upper-bound estimate of $\lambda_{\min}(H)$ without access to a reliable classical algorithm for estimating $\lambda_{\min}(H)$?
We address this question by developing a variational quantum algorithm (called dual-VQE) to estimate a lower bound on the ground-state energy, which can be compared to the upper-bound estimate obtained from VQE. Since the two estimates ideally sandwich the true optimal value, they serve as quality checks on each other. The key ideas of our approach are to 1) use SDP duality theory to reformulate $\lambda_{\min}(H)$ as a maximization problem and 2) further reformulate this maximization problem in such a way that it can be estimated by means of a variational quantum algorithm, which finally leads to a lower-bound estimate of $\lambda_{\min}(H)$. We call our method dual-VQE because it makes use of the dual characterization of $\lambda_{\min}(H)$ to arrive at this lower-bound estimate. After executing VQE and dual-VQE, one can then compare the upper-bound and lower-bound estimates to understand how well a quantum computer can approximate the true optimal value $\lambda_{\min}(H)$, without the aid of precise classical methods.
The dual-VQE method involves 1) converting inequality constraints to equality constraints by means of positive semi-definite slack variables, 2) converting a constrained optimization into an unconstrained one by means of a penalty term in the objective function, 3) replacing positive semi-definite slack variables with scaled parameterized quantum states that are efficiently preparable on quantum computers, and 4) estimating various terms in the objective function by means of the destructive swap test [21] and sampling estimates of expectations of observables. Our companion paper [17] explores this general approach in a much broader context beyond the ground-state energy problem.
Although there have been various classical algorithms proposed for obtaining lower bounds on the ground-state energy [28, 29, 30, 8, 23], they all ultimately suffer from the exponential increase of the space on which $H$ acts, as the number of particles increases. Our approach is fundamentally different, involving a reformulation of the original optimization problem in terms of its SDP dual, and then bounding that quantity from below using a variational quantum algorithm. Our approach is also complementary to [9, Section VI-A], which does not take a variational approach, nor does it provably obtain a lower bound on the ground-state energy.
The rest of our paper proceeds by detailing the dual-VQE method outlined above, and then we report the results of simulations of the transverse-field Ising model, a key example of physical interest. We find that the dual-VQE approach on a two-qubit example takes more iterations to converge as compared to the standard VQE optimization; however, it approaches the true value with error on the order of $10^{-2}$. We also note that in this example, the objective function values across training are noisier for dual-VQE when compared to VQE. We expect this behavior due to the increase in number of terms and parameters when estimating the ground-state energy with dual-VQE. Details of numerical simulations are given below, and training plots are shown in Figure 1.
Dual-VQE theoretical derivation—Let us first use some basic SDP duality theory to derive a well known result, that the following weak-duality inequality holds:
$\sup_{\eta\in\mathbb{R}}\left\{\eta:\eta I\leq H\right\}\leq\inf_{\rho\in\mathcal{D}}\operatorname{Tr}[H\rho],$ (4)
where $\sup_{\eta\in\mathbb{R}}\left\{\eta:\eta I\leq H\right\}$ is the SDP dual of (1). Note that actually equality holds in (4) due to Slater’s theorem [13]: we can pick $\rho=I/d$, where $d$ is the dimension of the underlying Hilbert space, and $\eta<\lambda_{\min}(H)$ as strictly feasible choices in both the dual and primal, thus satisfying the conditions of Slater’s theorem. Recall that the matrix inequality $A\geq B$ holds for Hermitian matrices $A$ and $B$ if $A-B$ is a positive semi-definite matrix. Then consider that
$\inf_{\rho\in\mathcal{D}}\operatorname{Tr}[H\rho]$ $=\inf_{\rho\geq 0}\left\{\operatorname{Tr}[H\rho]:\operatorname{Tr}[\rho]=1\right\}$ (5)
$=\inf_{\rho\geq 0}\sup_{\eta\in\mathbb{R}}\left\{\operatorname{Tr}[H\rho]+\eta\left(1-\operatorname{Tr}[\rho]\right)\right\}$ (6)
$=\inf_{\rho\geq 0}\sup_{\eta\in\mathbb{R}}\left\{\eta+\operatorname{Tr}[\left(H-\eta I\right)\rho]\right\}$ (7)
$\geq\sup_{\eta\in\mathbb{R}}\inf_{\rho\geq 0}\left\{\eta+\operatorname{Tr}[\left(H-\eta I\right)\rho]\right\}$ (8)
$=\sup_{\eta\in\mathbb{R}}\left\{\eta:\eta I\leq H\right\}.$ (9)
The second equality follows by introducing the Lagrange multiplier $\eta$ and because the constraint $\operatorname{Tr}[\rho]=1$ does not hold if and only if $\sup_{\eta\in\mathbb{R}}\left\{\eta\left(1-\operatorname{Tr}[\rho]\right)\right\}=+\infty$. The third equality holds by simple algebra. The sole inequality follows from the standard max-min inequality. The final equality follows by thinking of $\rho$ as a matrix Lagrange multiplier and because the constraint $\eta I\leq H$ does not hold if and only if $\inf_{\rho\geq 0}\left\{\operatorname{Tr}[\left(H-\eta I\right)\rho]\right\}=-\infty$.
Now that we have recalled this standard result, let us focus on showing how to estimate the lower bound in (4) on a quantum computer. Recalling that the matrix inequality $\eta I\leq H$ is equivalent to the existence of a positive semi-definite matrix $W$ such that $H-\eta I=W$, we conclude that
$\sup_{\eta\in\mathbb{R}}\left\{\eta:\eta I\leq H\right\}=\sup_{\eta\in\mathbb{R},W\geq 0}\left\{\eta:H-\eta I=W\right\}.$ (10)
The matrix $W$ is known in the theory of optimization [13] as a slack variable because it eliminates the slack in the inequality constraint $\eta I\leq H$ by reducing it to an equality constraint.
Our next observation is that every positive semi-definite matrix $W\neq 0$ can be written as a scaled density matrix, i.e., $W=\nu\omega$, where $\nu=\operatorname{Tr}[W]$ and $\omega=W/\nu\in\mathcal{D}$, implying that
$\sup_{\begin{subarray}{c}\eta\in\mathbb{R},\\ W\geq 0\end{subarray}}\left\{\eta:H-\eta I=W\right\}=\sup_{\begin{subarray}{c}\eta\in\mathbb{R},\nu\geq 0,\\ \omega\in\mathcal{D}\end{subarray}}\left\{\eta:H-\eta I=\nu\omega\right\}.$ (11)
Next, we adopt a standard approach of introducing a penalty term in the objective function to transform the constrained optimization to an unconstrained one:
$\sup_{\begin{subarray}{c}\eta\in\mathbb{R},\nu\geq 0,\\ \omega\in\mathcal{D}\end{subarray}}\left\{\eta:H-\eta I=\nu\omega\right\}\\ =\lim_{c\to\infty}\sup_{\begin{subarray}{c}\eta\in\mathbb{R},\nu\geq 0,\\ \omega\in\mathcal{D}\end{subarray}}\left\{\eta-c\left\|H-\eta I-\nu\omega\right\|_{2}^{2}\right\},$ (12)
where $c>0$ is a penalty parameter and $\left\|Z\right\|_{2}\coloneqq\sqrt{\operatorname{Tr}[Z^{{\dagger}}Z]}$ is the Hilbert–Schmidt norm of a matrix $Z$. Note that the equality in (12) holds whenever the penalty term is faithful (specifically, we invoke [6, Proposition 5.2.1] to justify this conclusion, setting $\lambda_{k}=0$ therein for all $k$). By faithfulness, we mean that the penalty term is equal to zero if and only if $H-\eta I=\nu\omega$. This is true for the Hilbert–Schmidt norm, and this particular norm has the convenient advantage that it can be estimated by quantum sampling procedures that we recall below.
The final step is to modify the optimization in (12) to be over a parameterized set $\left\{\omega(\theta)\right\}_{\theta\in\Theta}$ of density matrices instead of being over all density matrices, which implies that
$\sup_{\begin{subarray}{c}\eta\in\mathbb{R},\nu\geq 0,\\ \theta\in\Theta\end{subarray}}\left\{\eta-c\left\|H-\eta I-\nu\omega(\theta)\right\|_{2}^{2}\right\}\\ \leq\sup_{\begin{subarray}{c}\eta\in\mathbb{R},\nu\geq 0,\\ \omega\in\mathcal{D}\end{subarray}}\left\{\eta-c\left\|H-\eta I-\nu\omega\right\|_{2}^{2}\right\}.$ (13)
Piecing together all these steps in (4)–(13), expanding the Hilbert-Schmidt norm, and defining
$f^{\ast}(c)\coloneqq\sup_{\eta\in\mathbb{R},\nu\geq 0,\theta\in\Theta}f(\eta,\nu,\theta,c),$ (14)
where $\Theta$ is a parameter set,
$f(\eta,\nu,\theta,c)\coloneqq\eta-c\big{(}\operatorname{Tr}[H^{2}]+\eta^{2}2^{n}+\nu^{2}\operatorname{Tr}[\omega(\theta)^{2}]\\ -2\eta\operatorname{Tr}[H]-2\nu\operatorname{Tr}[H\omega(\theta)]+2\eta\nu\big{)}\,,$ (15)
and $c$ is a penalty parameter, we obtain the main theoretical result supporting the dual-VQE approach:
Theorem 1 (Dual-VQE lower bound)
The function $f^{\ast}(c)$ in (14) provides a lower bound on the ground-state energy of the Hamiltonian $H$ in the limit $c\to\infty$; i.e.,
$\lim_{c\to\infty}f^{\ast}(c)\leq\lambda_{\min}(H).$ (16)
Inspecting (15), only two terms $\operatorname{Tr}[\omega(\theta)^{2}]$ and $\operatorname{Tr}[H\omega(\theta)]$ depend on the optimization variable $\omega(\theta)$, while constant terms like $\operatorname{Tr}[H^{2}]$ and $\operatorname{Tr}[H]$ (having no dependence on $\eta$, $\nu$, or $\omega(\theta)$) can be calculated offline.
A key case of interest for physics and chemistry is when $H$ is a linear combination of Pauli strings, so that
$H=\sum_{\overrightarrow{x}}\alpha_{\overrightarrow{x}}\sigma_{\overrightarrow{x}},$ (17)
where $\overrightarrow{x}\in\left\{0,1,2,3\right\}^{n}$, $\alpha_{\overrightarrow{x}}\in\mathbb{R}$, $\sigma_{\overrightarrow{x}}\equiv\sigma_{x_{1}}\otimes\cdots\otimes\sigma_{x_{n}}$ is a Pauli string, and $\sigma_{0}\equiv I$, $\sigma_{1}\equiv\sigma_{X}$, $\sigma_{2}\equiv\sigma_{Y}$, and $\sigma_{3}\equiv\sigma_{Z}$ are the standard Pauli matrices. For this case, it follows from substitution and the orthogonality relation $\operatorname{Tr}[\sigma_{\overrightarrow{x}}\sigma_{\overrightarrow{y}}]=2^{n}\delta_{\overrightarrow{x},\overrightarrow{y}}$ that
$\operatorname{Tr}[H^{2}]$ $=2^{n}\left\|\overrightarrow{\alpha}\right\|_{2}^{2},$ (18)
$\operatorname{Tr}[H]$ $=2^{n}\alpha_{\overrightarrow{0}},$ (19)
$\operatorname{Tr}[H\omega(\theta)]$ $=\sum_{\overrightarrow{x}}\alpha_{\overrightarrow{x}}\operatorname{Tr}[\sigma_{\overrightarrow{x}}\omega(\theta)].$ (20)
[…]
Dual-VQE algorithm for lower bounding the ground-state energy—The basic idea behind the dual-VQE algorithm is to follow the standard penalty method [6, Section 5.2]: define a sequence $\left(f^{\ast}(c_{k})\right)_{k\in\mathbb{N}}$ of optimization problems where $\left(c_{k}\right)_{k\in\mathbb{N}}$ is a monotone increasing sequence of penalty parameters. We then solve the optimization for each $k$ and use the solution as the initial guess for the next iteration. Under the assumption that each optimization $f^{\ast}(c_{k})$ can be solved, the method is guaranteed to converge to the correct solution. In our simulations, we used a constant value of the penalty parameter $c$ and found that it performs well in practice.
In dual-VQE, we estimate the expectation $\operatorname{Tr}[H\omega(\theta)]$ by sampling from a quantum computer, and we use other approaches, either quantum or classical, for estimating the purity $\operatorname{Tr}[\left(\omega(\theta)\right)^{2}]$. If employing gradient descent for solving each optimization, one should estimate the gradient as well, which we discuss further below. The main additional overhead compared to VQE is the need to employ an ansatz for generating a mixed state $\omega(\theta)$ and to estimate the purity $\operatorname{Tr}[\left(\omega(\theta)\right)^{2}]$, given that VQE already estimates a term like $\operatorname{Tr}[H\omega(\theta)]$. However, as we discuss below, one approach reduces the estimation of the purity $\operatorname{Tr}[\left(\omega(\theta)\right)^{2}]$ to a classical sampling task, thus removing the need for a quantum computer for this part.
There are at least two general ways of parameterizing the set of mixed states, called the purification ansatz [16, 35, 19] and the convex combination ansatz [46, 26, 19, 41]. Focusing on the first one to start, it is based on the purification principle [12, 44, 45], in which every mixed state is realized as the marginal of a pure state on a larger system. For the purification ansatz, we simply start with a parameterized family $\left\{|\psi(\theta)\rangle_{RS}\right\}_{\theta\in\Theta}$ of pure states of a reference system $R$ and the system $S$ of interest, and then we obtain a parameterized family $\left\{\omega(\theta)\right\}_{\theta\in\Theta}$ of mixed states via the partial trace $\omega(\theta)=\operatorname{Tr}_{R}[|\psi(\theta)\rangle\!\langle\psi(\theta)|_{RS}]$. We can estimate the purity $\operatorname{Tr}[\left(\omega(\theta)\right)^{2}]$ by means of the destructive swap test [21] (see, e.g., [11, Section 2.2] for a precise statement of the algorithm). We can estimate the gradient of the objective function in (16) by means of the parameter-shift rule [27, 33, 39], simply because
$\operatorname{Tr}[H\omega(\theta)]=\langle\psi(\theta)|_{RS}\left(I_{R}\otimes H_{S}\right)|\psi(\theta)\rangle_{RS},$ (21)
and one can also use a slightly modified version of the parameter-shift rule along with the destructive swap test to estimate the gradient of the purity term $\operatorname{Tr}[\omega(\theta)^{2}]$ because
$\operatorname{Tr}[\omega(\theta)^{2}]=\langle\phi(\theta)|\left(I_{R_{1}R_{2}}\otimes F_{S_{1}S_{2}}\right)|\phi(\theta)\rangle,$ (22)
where
$|\phi(\theta)\rangle_{R_{1}R_{2}S_{1}S_{2}}\equiv|\psi(\theta)\rangle_{R_{1}S_{1}}\otimes|\psi(\theta)\rangle_{R_{2}S_{2}},$ (23)
$F$ is the unitary and Hermitian swap operator, and we have suppressed system labels in (22) for brevity (see also [19, Eq. (C2)] in this context).
The convex combination ansatz is based on the fact that every mixed state can be written as a convex combination of pure orthogonal states. Following this idea, we obtain a parameterized family $\left\{\omega(\varphi,\gamma)\right\}_{\varphi,\gamma\in\Theta}$ by separately parameterizing the eigenvalues and eigenvectors as follows:
$\omega(\varphi,\gamma)=\sum_{x}p_{\varphi}(x)U(\gamma)|x\rangle\!\langle x|U(\gamma)^{{\dagger}},$ (24)
where $\left(p_{\varphi}(x)\right)_{x}$ is a parameterized probability distribution, $U(\gamma)$ is a parameterized unitary, and $\left\{|x\rangle\right\}_{x}$ is the standard computational basis. The parameterized probability distribution $p_{\varphi}$ can be realized by a neural-network-based generative model [2, 10, 34] or a quantum circuit Born machine [7], while the parameterized unitary $U(\gamma)$ can be realized by a parameterized circuit.
Interestingly, although it is necessary to use a quantum computer to estimate the purity $\operatorname{Tr}[\left(\omega(\theta)\right)^{2}]$ for the purification ansatz, it is not necessary to do so when using the convex combination ansatz in conjunction with a neural-network-based generative model: it suffices to employ a classical sampling approach, thus further reducing the quantum resource requirements for dual-VQE (see also [19, Eqs. (20)–(23)] in this context). This is because $\operatorname{Tr}[\left(\omega(\varphi,\gamma)\right)^{2}]=\sum_{x}p_{\varphi}^{2}(x)$, so that one can repeatedly sample from $p_{\varphi}$ and perform a collision test to estimate $\sum_{x}p_{\varphi}^{2}(x)$. Thus, for the convex combination ansatz used in the aforementioned way, one only needs a quantum computer to estimate the expectation
$\operatorname{Tr}[H\omega(\varphi,\gamma)]=\sum_{x}p_{\varphi}(x)\langle x|U(\gamma)^{{\dagger}}HU(\gamma)|x\rangle,$ (25)
[…]
Purification ansatz Convex-combination ansatz
Objective function Refer to caption Refer to caption
Error Refer to caption Refer to caption
Penalty value Refer to caption Refer to caption
Figure 1: Convergence of dual-VQE and comparison to standard VQE for a two-qubit example problem instance. The solid line shows the median value of the estimate, the shaded region represents the interquartile range, while the ground truth is marked by the dashed line. We were able to achieve an error of $10^{-2}$ after $20,000$ iterations of training.
Results of numerical simulations—We now present the results of simulations that test the performance of dual-VQE on the transverse field Ising model, described by the following Hamiltonian of $n$ qubits:
$H_{\text{TFI}}=\sum_{i=1}^{n-1}J_{i}\sigma_{Z}^{i}\otimes\sigma_{Z}^{i+1}+\sum_{i=1}^{n}g_{i}\sigma_{X}^{i},$ (26)
where tensor products with identities are left implicit, $J_{i}$ is an interaction strength, and $g_{i}$ is related to the strength of the external field. In particular, we showcase our algorithm on the following two-qubit example problem instance:
$H=\sigma_{Z}\otimes\sigma_{Z}+\sigma_{X}\otimes\sigma_{I}+\sigma_{I}\otimes\sigma_{X}.$ (27)
Prior to training, we initialized the parameters $\eta$, $\mu$, and $\theta$. We set the parameters $\eta$ and $\nu$ in (12) to be equal to zero and one, respectively. We selected the components of the initial parameter vector $\theta$ for the state $\omega(\theta)$ uniformly at random from $[0,2\pi]$. The optimal state $\omega$ is almost always a full-rank mixed state. Thus, we used either a three-layer purification ansatz or a convex-combination ansatz with a two-layer quantum circuit Born machine and a two-layer unitary. The particular structure of the purification ansatz and the quantum circuit Born machine we use can be found in a figure in our companion paper [17].
In addition to setting the initializations of the parameters of our optimization, we assigned specific values to the hyperparameters. We chose the penalty parameter $c$ to have a constant value of $10$ throughout training, and we implemented a decreasing learning rate scheme. We initialized the learning rate to a maximum value of $0.1$, and every $100$ iterations, we checked whether the objective function value decreased three times in a row over the past $300$ iterations. In this case, we reduced the learning rate by a factor of two. We repeated this process until a minimum value of $0.01$ was met.
Training crucially depends on estimating the gradient, which allows us to pick the next set of parameters. We used the SPSA algorithm [42], which provides an unbiased estimate of the gradient at each step. We chose this method as the per-iteration cost of this algorithm is constant in the number of parameters, and we found that it works well in practice. If the norm of the gradient had a value greater than one, we normalized the gradient before performing the update step of gradient descent. We found that this prevents immediate divergence due to a potentially high initial penalty value. It also allowed us to use higher values of the learning rate than we would have been able to otherwise. Training concluded when an arbitrarily set maximum number of iterations was met.
The results of the simulations are displayed in Figure 1. In these plots, solid lines depict results of both the standard VQE algorithm (with fixed learning rate of $0.005$ and gradient estimated by normalized SPSA) and our dual-VQE approach. The solid lines show median values across ten independent runs, and the shading is the interquartile range. The dashed line is the true value of the ground-state energy estimated via the SDP. The variation displayed in the graphs comes from three sources of randomness: the randomized initializations of all parameterized quantum circuits, randomness in the SPSA algorithm for gradient estimation, and shot noise over $10^{12}$ shots for every measurement of a quantum circuit. In our simulations, we were able to achieve an error of order $10^{-2}$ after $20,000$ iterations of training. These results give evidence that using VQE and dual-VQE allows us to bound the true value from both above and below, respectively, as desired.
Discussion and conclusion—In conclusion, we proposed dual-VQE as a variational quantum algorithm for lower bounding the ground-state energy of a Hamiltonian. This lower bound can be compared with the upper bound obtained from the traditional VQE method, and the two bounds serve as quality checks on each other that ideally determine an interval containing the optimal ground-state energy. As discussed above, the quantum computational resources required for evaluating the objective function of dual-VQE are no greater than those needed for VQE’s objective function when using the convex combination ansatz in conjunction with a classical generative model.
While Theorem 1 guarantees that dual-VQE provides a rigorous lower bound on the ground state energy in the limit $c\rightarrow\infty$, for finite $c$ the value provided by dual-VQE is technically an estimate rather than a strict bound. However, our numerical simulations give evidence, for the example considered, that VQE and dual-VQE do indeed bound the optimal ground-state energy from above and below, respectively. Nonetheless, it would be valuable to investigate whether it is possible to obtain bounds on the degree to which deviations are possible for finite $c$. This would allow us to certify the reliability of dual-VQE and serve as a guide to an optimum choice for $c$.
A key issue that plagues variational quantum algorithms is the presence of barren plateaus in the training landscape [31, 20, 37], meaning that the gradient (or equivalently cost differences [1]) vanishes exponentially fast in the number of qubits. This has been shown to occur for highly expressive [24] or highly entangling [32] ansätze or even at low circuit depths for global costs [15]. For local problem Hamiltonians, the only term we need to evaluate on a quantum computer, $\text{Tr}[H\omega(\theta)]$, is local and hence in such cases, for $\log(n)$-depth hardware efficient ansätze, dual-VQE provably avoids barren plateaus [15]. However, a worry here is that such short-depth circuits might also be classically simulable via state-of-the-art methods [43, 4, 25, 3, 38]. An important question for dual-VQE, shared also with other variational quantum algorithms, is thus whether we can identify quantum models that both avoid barren plateaus and cannot be classically simulated. One potential avenue would be to try and find clever initialization strategies to explore small subsections of the landscape that exhibit substantial gradients and are also hard to simulate classically.
For future work, it would certainly be of interest to scale up dual-VQE to much larger systems beyond the reach of brute-force classical simulation and to compare the results of VQE and dual-VQE with each other. We also think it would be worthwhile to pursue other methods, besides the penalty method, for performing the optimization in dual-VQE. A promising approach would be to combine a variational approach with an interior-point method, but for this approach, it seems we would need an efficient method, suitable for near-term quantum computers, for estimating the logarithm of the determinant of a matrix on a quantum computer.
Data availability statement—All source codes used to run the simulations and generate the figures are available as arXiv ancillary files with the arXiv posting of our companion paper [17].
Acknowledgements.
We thank Paul Alsing, Ziv Goldfeld, Daniel Koch, Saahil Patel, and Manuel S. Rudolph for helpful discussions. JC and HW acknowledge support from the Engineering Learning Initiative in Cornell University’s College of Engineering. ZH acknowledges support from the Sandoz Family Foundation Monique de Meuron program for Academic Promotion. IL, TN, DP, SR, and MMW acknowledge support from the School of Electrical and Computer Engineering at Cornell University. TN, DP, SR, and MMW acknowledge support from the National Science Foundation under Grant No. 2315398. DP, SR, and MMW acknowledge support from AFRL under agreement no. FA8750-23-2-0031. This material is based on research sponsored by Air Force Research Laboratory under agreement number FA8750-23-2-0031. The U.S. Government is authorized to reproduced and distribute reprints for Governmental purposes notwithstanding any copyright notation thereon. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of Air Force Research Laboratory or the U.S. Government. This research was conducted with support from the Cornell University Center for Advanced Computing, which receives funding from Cornell University, the National Science Foundation, and members of its Partner Program.